The sum of two numbers is -15. Their product is -100. Find the two numbers.
Answer by checkley75(3666)

X+Y=-15 X=-Y-15
X*Y=-100 NOW REPLACE X BY (-Y-15)
(-Y-15)*Y=-100
-Y^2-15Y+100=0
Y^2+15Y-100=0
(Y+20)(Y-5)=0
Y+20=-0
Y=-20 ANSWER.
Y-5=0
Y=5 ANSWER.
SAME ANSWERS FOR X.